Attendees: Glen, Xinyu, Riku, Don, Ben C, Anton, Joanna, Matteo
- Glen: Plugins? TFC-121 MBed TLS plugins
- Riku: Still working on it. STG-2633
- Glen: Staffing.
Theo onboard, setting up LAVA at home and will be integrating next
Leonardo remains onboard for a bit more time, will continue on MBed
Arthur getting NXP boards into LAVA as well as TFC-82 Github auth.
support to LAVA - patch applied, need LAVA update now
- Glen: Recent LAVA change creating a Code Coverage regression. Email
thread only. TF-A was working, now not. So a new ticket will be created
- Glen: backlog / sc-approved cleanup review
In progress: Reviewed, staffed, and all looks good
SC Approved: Expect script to be started w/ Paul
Several others are queued up.
TFC-29 - Joanna to check to see if this needs worked on. Matteo, also
want to visit backlog to see if higher priorities. Moved back to backlog.
- Anton: TFC-33 and 135 in backlog may be really simple - permission
- Joanna: TFC-75 - would be good to start on this one, some sync issues
on gerrit. Hoping it's another quick one. Joanna to follow up and update
- Joanna to prioritize TF-A tickets in backlog then we'll prioritize
those against other backlog items
*Attendees*: Don, Joanna, Matteo, Xinyu, Riku, Janos Follath, Anton, Dave
Rodgman, Ben Copeland, Shebu
- Glen: Staffing - Updated team on the staffing updates
- Theo and Paul joining.
- Leonardo transitioning over the next few weeks.
- Glen to meet w/ whole team tomorrow
- Leonardo still best person to contact on MBedTSL
- M2 completed, M3 blocked on need for plugins
- Riku to take care of plugins, targeting today to complete
- Still completing
- TFC-82: Arthur was working on this. Want his updates released and
- TFC-87: Arthur has been working w/ Joanna on this. May take a couple
of passes, but initial solution in the next day
- TFC-183 - MUSCA B are off line.
- Ben: Dave looking into this. Failed due to debian update and debian
- Glen: Back it out?
- Ben recommends moving forward and resolving in next sprint (next
- Anton: Only MUSCA B offline? Preparing for next month release, so
need by then for sure.
- Agreed to put in next week
- Glen: NXP Boards coming up in Lab
- Glen: Chromebooks (Jacuzi - TFC-45) received in Cambridge, got out of
- Glen: Creating card for next set of Chromebooks (Lazor)
- Glen: Renesas beginning to ship the h/w
- Shebu: Renesas Intend to push board support for TF-M
- Matteo: Any priority calls to make?
- Glen: Some of the smaller backlog tickets. Glen will be moving items
out of backlog and assigning them.
- Matteo: To draft an email on backlog email and share w/ Glen.
- Shebu: Some SC Approved have yet to be touched.
- How much effort for Disaster Recovery? It could be a larger task.
Will break this down into subtasks and do the assessment.
- ACTION: Joanna, Anton, Matteo, Shebu to go thru backlog and SC
approved and make adjustments in the Kanban. Keep others updated via email.
Attendees: Don, Riku, Xinyu, Shebu, Dave Rodgman, Ben, Joanna
- Several platforms ready to add to the Cambridge lab. Includes NXP,
- TFC-181: patches in review
- TFC-176: Continue to work. Moed from SC Approved to In Progress.
- May have a temporary fix from Leonardo.
- Monitoring at this point. Joanna recommends holding off on ticket
TFC-96 until TFC-176 is stable so as not to add more variables into the
solution. After a few days of stability, can come back to this.
- Moved TFC-36 to REOPENED and prioritized
- Moved TFC-36 to SC Approved
- Verify MBed status. Still in Blocked.
- Action: Don check w/ Glen/Leonardo as to why the Mbed work still in
- Anything else to work on here?
- Joanna: TFC-87: Would be good to start if time allows. Doesn't have
to all be done at once.
- Shebu: Update on staffing
- Two engineerings from Linaro coming on to help for next 6 months to
supplement the staff. Theodore Grey & Paul Sokolovskyy (Welcome aboard!)
- Will want to make sure the team is appropriately loaded.
- Board additions will continue.
- MBedTLC CI up and running important too.
- Over the next couple of weeks Glen will be integrating them into
the team as they spin up.
*Attendees*: Don, Glen, Anton, Shebu, Matteo, Joanna, Ben C, Dave Rodgman,
- Servers went down - certificate expiration issue resolved. Blocker
- TFC-179/TFC-80 (duplicate) - certificates updated and these can be
- Glen: Adding boards:
- Have documentation on boards. Should we add vendor boards to
- User's manual.
- Determine if public?
- Don: Assumed it was public.
- Shebu: Worth asking Eric, Julius, Sean, and others if they're OK to
- Glen: Can share what would be published in the User's Guide first.
- Would be OK to point to vendor info if works
- Anton: TF-A/M docs could also point to this
- Platform integration updates:
- Glen: Cypress/NXP in Cambridge. Will get integrated shortly.
- Google Chromebooks.
- Arthur and Fathi's returned to sender, the 3 platforms made it
to the lab.
- Working with Google to start the 2nd platform.
- Dave's team provided the AMI files.
- Dave: Darryl expected additional AMI's beyond Ubuntu image. Will
build additional configs from those after Ubuntu.
- TFC-176: Leonardo has a work-around (3 retries), but not getting to
root cause. Likely need a longer-term solution.
- Glen asking for Expect script review from Joanna.
- TFC-82: Github auth. Working, now looking to deploy.
- Anton: TF-M health check job. Expected OpenCI team to keep it up.
Sanity test is failing.
- Don: Suggest creating a ticket.
- Action: Glen/Ben to open a ticket so that can track/assign.